OpenStudy (meilendurcer):

What three numbers come next in this number pattern? 1, 8, 27, 64, __, __, __, … A. 25, 36, 49 B. 125, 216, 343 C. 100, 121, 144 D. 75, 88,

OpenStudy (anonymous):

x^3 is the pattern, 125, 6^3, 7^3

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@ sms, Can you perhaps remind us what the formula is to determine lets say n = 45?

OpenStudy (anonymous):

45^3. I can't explain how I knew it was this pattern besides recognizing that 8 = 2^3 and 27= 3^3, and so on. n^3 is the formula

OpenStudy (anonymous):

Yes you are right and very bright to have noticed it too, but this is not always the situation for these rows. I remember there was an equation that one could use to determine the values of these rows. Just can't find the English term for it to look it up...